OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5? A. y = 5x - 3 B. y = 5x - 1 C. y = 5x + 3 D. y = 5x - 19

OpenStudy (anonymous):

if you use \[y - y_{1}=m(x-x{1})\] and plug in your slope and x and y and solve for y you will get a \[y = mx + b\] finsh and you will have your answer

OpenStudy (anonymous):

Or use y=mx+b where m is slope, and our slope is 5 and b is y intercept so plug in what we know y=5x+b and the like has to go through the point (4,1) so plug in 4 for the x and 1 for y and solve for b 1=5(4)+b b=-19 final answer: y=5x=19

OpenStudy (anonymous):

^ yep thank you condawg :)

OpenStudy (anonymous):

=) np

OpenStudy (anonymous):

Thank u that helped me to understand it better

OpenStudy (anonymous):

do you understand everything I said?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ok